OpenStudy (theedwardsfamily):

@zbratz7 @felicity214

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

distribute the 1/5 to the 2x and -15. You get (0.4x - 3)

OpenStudy (theedwardsfamily):

ok

OpenStudy (anonymous):

0.4x - 3 + 3/5x = 4 and 1/4

OpenStudy (anonymous):

I am going to switch the fractions to decimals because I think they are much easier.

OpenStudy (theedwardsfamily):

ok

OpenStudy (anonymous):

So when you have it as decimals you get: 0.4x - 3 + 0.6x = 4.25

OpenStudy (theedwardsfamily):

Yes I get.

OpenStudy (anonymous):

Add like terms to get: 0.4x + 0.6x -3 = 4.25 x -3 = 4.25 +3 +3 x = 7.25

OpenStudy (anonymous):

in fraction form: 7 and 1/4

OpenStudy (theedwardsfamily):

Whoop! I passed! Thnx @Felicity214 ! :D

OpenStudy (anonymous):

your welcome :) What did you pass with?

OpenStudy (theedwardsfamily):

80%

OpenStudy (theedwardsfamily):

Thanks again!😊